# CR3DT: Camera-RADAR Fusion for 3D Detection and Tracking

### v1

@@ V-D CC-3DT++ Tracking Architecture Ablation
To better understand the different impacts of our changes to the original CC-3DT [3] tracking architecture and the addition of our velocity-based affinity term, we conduct extensive ablation studies depicted in Table IV. Furthermore, we report the performance of the different tracking architecture configurations on both our baseline BEVDet detector as well as our own CR3DT detector in Table III to investigate the effects of our changes more generally.
@@
We ran three main experimental studies, the first was an investigation of the matching score threshold utilized in the greedy matching algorithm of CC-3DT, which is tightly coupled to the affinity scores calculated in Eq. 1. We observe significant performance gains both in an increase of AMOTA and a stark decrease in IDS for a slightly lower threshold, with a stark performance decrease for too low of a threshold, as indicated in Table III(a). Secondly, we explored different weightings of the embeddings correlation and motion correlation terms in Eq. 1. As shown in Table III(b), we find that a larger weight on the motion correlation term leads to an AMOTA gain, although at the cost of an IDS increase. Lastly, we combined the best threshold and weighting configuration and examined the motion correlation term in Eq. 2 itself. Trying to leverage the much-improved velocity estimations of our detector, we exchanged the original trade-off parameter (the cosine similarity between the predicted motion direction and the observed motion direction in the xy-plane) with a new velocity similarity term, which compares the predicted with the observed velocity directly as explained in Eq. 3. As hypothesized, we observe a further performance increase in both AMOTA and IDS in Table III(c).
It additionally has to be noted, that we fixed a minor bug in the original implementation of CC-3DT, leading to the performance increase observed in the default configurations shown in the ablations in comparison to Table III (32.3% mAP vs. 31.2% mAP) .

### v2

@@ V-D CC-3DT++ Tracking Architecture Ablation
To better understand the impact of our different changes to the original CC-3DT [3] tracking architecture and the addition of our velocity-based affinity term, we conduct extensive ablation studies depicted in Table IV. Furthermore, we report the performance of the different tracking architecture configurations on both our baseline BEVDet detector as well as our own CR3DT detector in Table III to investigate the effects of our changes more generally.
@@
We ran three main experimental studies, the first was an investigation of the matching score threshold utilized in the greedy matching algorithm of CC-3DT, which is tightly coupled to the affinity scores calculated in Eq. 1. We observe significant performance gains both in an increase of AMOTA and a stark decrease in IDS for a slightly lower threshold, with a stark performance decrease for too low of a threshold, as indicated in Table IV(a). Secondly, we explored different weightings of the embeddings correlation and motion correlation terms in Eq. 1. As shown in Table IV(b), we find that a larger weight on the motion correlation term leads to an AMOTA gain, although at the cost of an IDS increase. Lastly, we combined the best threshold and weighting configuration and examined the motion correlation term in Eq. 2 itself. Trying to leverage the much-improved velocity estimations of our detector, we exchanged the original trade-off parameter (the cosine similarity between the predicted motion direction and the observed motion direction in the xy-plane) with a new velocity similarity term, which compares the predicted with the observed velocity directly as explained in Eq. 3. As hypothesized, we observe a further performance increase in both AMOTA and IDS in Table IV(c).
It additionally has to be noted, that we fixed a minor bug in the original implementation of CC-3DT, leading to the performance increase observed in the default configurations shown in the ablations in comparison to Table III (32.3% AMOTA vs. 31.2% AMOTA) .